# Tell Me What You Don’t Know: Enhancing Refusal Capabilities of Role-Playing Agents via Representation Space Analysis and Editing

### v1

@@ 6.3.1 Evaluation on MT-Bench
  Method Llama-3.1-8B-Instruct Llama-3-8B-Instruct Mistral-7B-Instruct-v0.2
@@
Table 5: Results of evaluations on different models and methods for MT-Bench. MT-Bench contains 8 subtasks, MT-Bench-Roleplay is one of the subtasks. Representation Editing demonstrates good performance not only in roleplay but also in general conversation.
@@ 7 Conclusion
Our study investigated PRAs capabilities in handling conflicting requests, with a focus on enhancing their ability to recognize and refuse inappropriate queries. Our evaluation of state-of-the-art models revealed significant performance differences across different conflict scenarios, particularly in dealing with parametric knowledge conflicts. Through analysis of model representations, we uncovered the existence of distinct representation spaces for different roles and conflict types within the models. This key finding explains the observed performance differences and provides a foundation for targeted improvements in RPA design. Our proposed representation editing approach offers a promising solution for enhancing RPAs’ refusal capabilities without training.
Appendix A Evaluation Protocol
@@
Evaluation of General Conversation Ability. General conversation ability is the foundational capability of RPAs. Assessing the general conversation ability of role-playing models is crucial because it directly impacts the user experience and satisfaction during interactions with the model. General conversation ability includes consistency, quality, and factuality, which collectively determine the fluency, depth, and accuracy of the conversation (Mesgar et al., 2020; Zhang et al., 2021; Tu et al., 2024).
@@
To assess RPAs’ performance across these dimensions, we use GPT-4o to score them. The specific scoring criteria for each dimension can be found in Appendix B
Appendix B Prompt
@@
All prompts we used are listed at Figures 5, 6, 7, 8. For evaluation, we listed our scoring criteria in Table 6
Figure 5: Role-playing Promp
@@ 9. Factuality of Response
Appendix C Training Details
C.1 Fine-tuning Details
For supervised fine-tuning and LoRA, we used the following experimental setup and hyperparameters:
  • Precision: Float32
  • Epochs: 1
  • Weight Decay: 0
  • Warmup ratio: 0.03
  • Learning rate: $2e^{-5}$
  • Max Seq. length: 2,048
  • Effective batch size: 128
For LoRA training, we used the following:
  • Precision: Float32
  • Epochs: 1
  • Weight Decay: 0
  • Warmup ratio: 0.03
  • Learning rate: $3e^{-4}$
  • Learning rate scheduler: cosine
  • Max Seq. length: 2,048
  • Effective batch size: 128
  • Lora rank: 16
  • Lora alpha: 16
  • Lora dropout: 0.1
C.2 Linear Probe Details
We trained linear probes using 200 queries each for non-conflict, role setting conflict, and factual knowledge conflict scenarios. For evaluation, For the five types of queries, we selected 50 samples each and used the average accuracy of role setting conflict and role profile conflict as the accuracy for contextual conflict. Similarly, we used the average accuracy of factual knowledge conflict and absent knowledge conflict as the accuracy for parametric knowledge conflict.
For the training of linear probes, we used 6 different random seeds. The experimental setup and hyperparameters for the linear probes are as follows:
  • Training epochs: 10
  • Learning rate: $5e^{-5}$
  • Learning rate scheduler: linear
  • Network structure: 3-layer fully connected network with dimensions (model_hidden_state, 512, 2)
  • Optimizer: Adam optimizer
Appendix D More analysis
D.1 More Analysis of Probe Result
From the Figure 2 we can also observe the following phenomenon:
In order to verify the above phenomenon, we apply the representation of the refusal direction obtained from Llama3.1-8B-Instruct to Llama3-8B-Instruct, as shown in Table 7.
  Non-conflict Role Setting Role Profile Factual Knowledge Absent Knowledge Average
Llama3-8B-Instruct $1.88$ $1.94$ $1.62$ $1.03$ $0.75$ $1.44$
w/ Llama3-8B-Instruct rejection direction $1.87$ $1.96$ $1.69$ $1.17$ $0.89$ $1.52$
w/ Llama3.1-8B-Instruct rejection direction $1.87$ $1.96$ $1.71$ $1.17$ $0.92$ $1.53$
Table 7: Model feature similarity verification experiment
D.2 Analysis of Representation Editing Method
Well performance in contextual conflicts In the two conflict types directly related to the character - “Role Setting” and “Role Profile” - the representation editing method demonstrated excellent performance across all models, typically outperforming or matching other methods.
Improvement in parametric knowledge conflicts In the two conflict types involving parametric knowledge - “Fact Knowledge” and “Absent Knowledge” - the representation editing method significantly outperformed FT and LoRA methods in most cases. This improvement is particularly evident in the Llama3-8B-Instruct and Mistral-7B-Instruct models.
D.3 Analysis of Representation via t-SNE

### v2

@@ 5.3.1 Main Evaluation Result
To avoid potential bias from using GPT-4o for both data generation and evaluation, we report results using different LLMs as evaluators in Table 6.
5.3.2 Human Evaluation Result
To validate our automated evaluation result and further assess the effectiveness of different methods, we conducted a human evaluation study. We recruited five novel enthusiasts to evaluate Llama-3-8B-Instruct outputs. For each query type of each role, we randomly sampled 10 examples for assessment. The evaluators followed the same nine-dimensional scoring criteria used in our automated evaluation, ensuring consistency in the assessment framework. The results, presented in Table 5, demonstrate strong alignment with our automated evaluation findings. The Representation Editing method achieved comparable or better performance acr0oss different query types. This human evaluation validates that our approach effectively enhances the model’s refusal capabilities without compromising its general role-playing abilities.
@@ 5.3.3 Evaluation on MT-Bench
  Method Llama-3.1 Llama-3 Mistral
@@
Table 7: Results of evaluations on different models and methods for MT-Bench. MT-Bench contains 8 subtasks, MT-Bench-Roleplay is one of the subtasks. The model parameters are $7$B or $8$B.Representation Editing demonstrates good performance not only in roleplay but also in general conversation.
@@ 6 Conclusion
Our study investigated RPAs capabilities in handling conflicting requests, with a focus on enhancing their ability to recognize and refuse inappropriate queries. Our evaluation of state-of-the-art models revealed significant performance differences across different conflict scenarios, particularly in dealing with parametric knowledge conflicts. Through analysis of model representations, we uncovered the existence of distinct representation spaces for different roles and conflict types within the models. This key finding explains the observed performance differences and provides a foundation for targeted improvements in RPA design. Our proposed representation editing approach offers a promising solution for enhancing RPAs’ refusal capabilities without training.
Limitations
While our study demonstrates the effectiveness of representation editing for enhancing refusal capabilities in models with 7-8B parameters, extending this approach to larger state-of-the-art models (such as those with 70B+ parameters) represents an important direction for future research. As model scale increases, the complexity of representation spaces and interaction patterns may present new challenges for our editing method. Additionally, while we focused on role-playing scenarios in role kownledge QA, exploring the applicability of representation editing in other domains could reveal new insights about the method’s generalizability.
Reproducibility Statement
We have publicly shared our code and dataset through a GitHub repository https://github.com/LiuAmber/RoleRef. To further ensure replicability, we asked a colleague unfamiliar with our method to install and test. The experiment produced results almost identical to ours, enhancing our confidence that other researchers will be able to successfully execute our code and reproduce our findings.
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments. This work was supported by National Natural Science Foundation of China (No. 62076068).
Appendix A Related Work
A.1 Role-Playing Agents
Enhancing Role-Playing Performance. Methods to improve RPAs are broadly categorized into prompt-based and fine-tuning-based approaches. Prompt-based methods provide models with detailed character descriptions, outlining attributes such as age, personality, and abilities, to facilitate accurate role-playing (Wang et al., 2024a; Zhou et al., 2024). Fine-tuning-based methods involve training models on role-specific behaviors, often using data sourced from manual annotations (Zhou et al., 2024; Chen et al., 2023; Zhang et al., 2024b), online resources (Zheng et al., 2019; Qian et al., 2021; Song et al., 2020; Shao et al., 2023; Tu et al., 2024), or generated by LLMs (Wang et al., 2024a; Li et al., 2023; Zhao et al., 2024a; Ahn et al., 2024; Lu et al., 2024). These methods aim to instill role-consistent behaviors and dialogue patterns in the models.
A.2 Knowledge Boundaries and Refusal Strategies
Parameteric Knowledge. Yang et al. (2023) and Cheng et al. (2024) explore teaching models to express uncertainty using prompt-based, fine-tuning, and preference-aware optimization methods. Xu et al. (2024a) propose a reinforcement learning method based on knowledge feedback to dynamically determine the model’s knowledge boundaries. Similarly, Zhang et al. (2024a) identifies knowledge gaps between pre-trained parameters and instruction-tuning data, constructing refusal-aware data by appending uncertainty expressions and improving the model’s ability to answer known questions while refusing unknown ones. Chen et al. (2024c) detect the knowledge boundaries of LLMs through internal confidence and teach LLMs to recognize and express these boundaries. Zhao et al. (2024c) propose a self-detection scheme to identify unknown knowledge by examining behavioral differences under varying formulations and the atypicality of input expressions. To address factual errors and outdated knowledge in parameterized knowledge, mainstream methods convert parameterized knowledge into contextual knowledge.
Contextual Knowledge. Cao (2024) use an independent structured knowledge base to represent the knowledge scope of LLMs, making LLMs process input-output data without relying on internal knowledge, thereby avoiding misinformation. Prompting LLMs to refuse to answer difficult questions improves system reliability. Deng et al. (2024) generate extensive unknown question-response data through class-aware self-augmentation and select qualified data via differential-driven self-curation, fine-tuning LLMs to improve their response capabilities to various unknown questions, enabling the model to refuse and explain why it cannot answer. Brahman et al. (2024) categorize scenarios requiring refusal to answer, and explore different training strategies to teach models to say “no." Zhao et al. (2024b) investigate decision boundaries in in-context learning by analyzing decision boundaries in binary classification tasks.
Appendix B Evaluation Protocol
@@
Evaluation of General Conversation Ability. General conversation ability is the foundational capability of RPAs. Assessing the general conversation ability of role-playing models is crucial because it directly impacts the user experience and satisfaction during interactions with the model. General conversation ability includes consistency, quality, and factuality, which collectively determine the fluency, depth, and accuracy of the conversation (Mesgar et al., 2021; Zhang et al., 2021; Tu et al., 2024).
@@
Figure 5: Screenshots of our human evaluation interface for rating. In each instance, evaluators scores according to the evaluation dimensions.
@@
To assess RPAs’ performance across these dimensions, we use GPT-4o to score them. The specific scoring criteria for each dimension can be found in Appendix E.For Human Evaluation, Figure 5 shows a screenshot of the interface used for our evaluation, which all evaluators utilized to rate the data.
Appendix C Details
C.1 Baselines Details
For supervised fine-tuning, we used the table 8 experimental setup and hyperparameters:
  Hyperparameter Value
  Precision Float32
Epochs $1$
Weight Decay 00
Warmup ratio $0.03$
Learning rate $2e^{-5}$
Max Seq. length $2048$
Effective batch size $128$
Table 8: Experimental Setup and Hyperparameters for Supervised FT
For LoRA training, we used table 9 experimental setup and hyperparameters:
  Hyperparameter Value
  Precision Float32
Epochs $1$
Weight Decay 00
Warmup ratio $0.03$
Learning rate $3e^{-4}$
Learning rate scheduler cosine
Max Seq. length $2048$
Effective batch size $128$
Lora rank $16$
Lora alpha $16$
Lora dropout $0.1$
Table 9: Experimental Setup and Hyperparameters for LoRA
C.2 Linear Probe Details
  1. 1.
    Data Preparation:
    • Hidden Representation Extraction: For each query, we first use the prompt shown in Figure 5 as input to the model. During the model’s forward pass, we extract the hidden states from a specified layer (e.g., the penultimate layer) to use as feature vectors.
    • Dataset Construction: We collect the corresponding hidden representations for different types of queries:
      • Training: 200 samples each for non-conflict, role setting conflict, and factual knowledge conflict scenarios
      • Testing: 50 samples for each of the five query types
      • For contextual conflict accuracy: average of role setting conflict and role profile conflict accuracies
      • For parametric knowledge conflict accuracy: average of factual knowledge conflict and absent knowledge conflict accuracies
    • Label Assignment: For binary classification, we assign a label of 1 to non-conflict query samples and a label of 0 to conflict query samples.
  2. 2.
    Model Definition:
    • Linear Probe Structure: We use a 3-layer fully connected network with dimensions (model_hidden_state, 512, 2) and an output layer with a Sigmoid activation function. This setup is used to probe whether the model perceives a query as conflicting with its knowledge.
  3. 3.
    Training Process:
    • Loss Function: We use the Mean Squared Error Loss (MSELoss) to optimize the model parameters.
    • Optimizer and Hyperparameters:
      • Optimizer: Adam optimizer
      • Learning rate: $5e^{-5}$
      • Learning rate scheduler: linear
      • Batch size: 512
      • Training epochs: 10
    • Training Strategy: The model is trained on the training set, and at the end of each epoch, its performance is evaluated on the validation set. The model parameters with the highest validation accuracy are saved.
  4. 4.
    Result Evaluation:
    • Evaluation Metrics: We calculate the prediction accuracy for each query type on the test set to assess the linear probe’s performance in distinguishing between different types of queries.
    • Experiment Reproducibility: To ensure the reliability of the results, we use 6 different random seeds and conduct experiments on data from multiple roles, calculating the average performance.
Figure 6: Methods to improve the model’s ability to refuse to answer.
C.3 Definitions of Refusal and Direct Response Region
  • Rejection Regions: When the similarity between the input query’s representation vector $\mathbf{h}^{l}$ and the rejection direction vector $\mathbf{d}^{\prime l}$ exceeds a certain threshold $\theta$, i.e., $\text{sim}(\mathbf{h}^{l},\mathbf{d}^{\prime l})\geq\theta$, the model is more inclined to trigger the refusal mechanism and decline to answer the query.
  • Direct Response Regions: When the similarity is below the threshold $\theta$, i.e., $\text{sim}(\mathbf{h}^{l},\mathbf{d}^{\prime l})<\theta$, the model tends to generate a direct response to the query.
Appendix D More analysis
D.1 Validating Results with Different LLM Evaluators
To validate the robustness of our evaluation framework, we assessed model performance using three different state-of-the-art LLMs as evaluators: qwen-max-2025-01-25, gemini-2.0-pro-exp-02-05, and Doubao-1.5-pro-32k-250115. Table 6 presents the evaluation results across different methods and query types.
The results demonstrate consistent patterns across all evaluator models. The Representation Editing method maintains competitive performance, achieving an average score of 1.54 across all evaluators, comparable to FT (1.55) and LoRA (1.55). This consistency is particularly evident in handling non-conflict queries (1.94) and contextual knowledge conflicts (1.97 for Role Setting), where the method performs strongly regardless of the evaluator model.
Notably, all evaluator models identify similar performance patterns across different query types. They consistently show that models perform better on contextual knowledge conflicts compared to parametric knowledge conflicts, aligning with our main findings using GPT-4o as the evaluator. This cross-model validation strengthens the reliability of our evaluation framework and the effectiveness of our proposed method.
D.2 More Analysis of Probe Result
From the Figure 3 we can also observe the following phenomenon:
In order to verify the above phenomenon, we apply the representation of the refusal direction obtained from Llama3.1-8B-Instruct to Llama3-8B-Instruct, as shown in Figure 7.
Figure 7: Model feature similarity verification experiment.
D.3 Analysis of Representation Editing Method
Well performance in contextual conflicts In the two conflict types directly related to the character - “Role Setting" and “Role Profile" - the representation editing method demonstrated excellent performance across all models, typically outperforming or matching other methods.
Improvement in parametric knowledge conflicts In the two conflict types involving parametric knowledge - “Fact Knowledge" and “Absent Knowledge" - the representation editing method significantly outperformed FT and LoRA methods in most cases. This improvement is particularly evident in the Llama3-8B-Instruct and Mistral-7B-Instruct models.
D.4 Analysis of Representation via t-SNE
D.5 Analysis of Computation Overhead
The representation editing method does not incur significant additional computational overhead. We analyze the computational overhead of our method mainly from two aspects: training overhead and inference overhead.
1. Training Overhead: As we have shown in Table 4 of our paper, our method does not involve any trainable parameters. Specifically, we only need to precompute and store the rejection vectors, which can then be simply added to the model’s internal representations during practical applications. Therefore, compared to FT and LoRA, the computational overhead during the training phase of the representation editing method is nearly zero.
2. Inference Overhead:During inference, our method only requires a simple vector addition operation between the precomputed rejection vectors and the current internal representations of the model. This operation has a computational complexity similar to the adapter modules in LoRA. Since this operation is extremely lightweight, its impact on inference time and computational resources is almost negligible. Therefore, our method does not introduce significant additional overhead during the inference phase either.
Appendix E Prompt
@@
All prompts we used are listed at Figures 10, 11, 12, 13. For evaluation, we listed our scoring criteria in Table 10
Figure 10: Role-playing Prompt